OpenStudy (anonymous):

Two particles are moving on an axis. The positions for the two particles when t is less than or equal to 0 are: s1(t)=t^2-5t-50 and s2(t)= t^3/3-4t^2-9t. When are the particles traveling at the same velocity? (I already have the first and second derivative for both functions)

OpenStudy (ivancsc1996):

Well if they travel at the same velocity then the first derivative of both should be the same :)

OpenStudy (agent0smith):

Set the first derivatives equal to each other, then solve for t.

OpenStudy (anonymous):

i thought i would have to set the derivatives equal to eachother and solve for t which what i began to do until i got stuck

OpenStudy (ivancsc1996):

I am sorry I was wrong.

OpenStudy (ivancsc1996):

What you have to do is find the first derivative of both and solve for t. I'll show you the maths.

OpenStudy (agent0smith):

Show us where you got stuck, @jstephanie \[\large v _{1}(t)=2t-5 \] \[\large v _{2}(t) = t^2 -8t -9\] Equate them to find where they're equal: \[\large 2t-5 = t^2 -8t -9\]

OpenStudy (ivancsc1996):

\[\frac{ d }{ dt }S _{1}=\frac{ d }{ dt }S _{2}=v _{1}=v _{2}=2t-5=t ^{2}-8t-9 \rightarrow 0=t ^{2}-10t-4\] Then you apply the quadratic formula and:\[t=\frac{ 10\pm \sqrt{10^{2}-4(-4)} }{ 2 }=\frac{ 10\pm \sqrt{116} }{ 2 }\]

OpenStudy (anonymous):

I didnt set it up like that thats probably where I went wrong @ivancsc1996 I did this: \[2t-5=t ^{2}-8t-9\rightarrow 2t= t ^{2}-8t-4\rightarrow 0= t ^{2}-6t-4\] then i got stuck because i couldnt get it to be a binomial

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

You made some mistakes, @jstephanie when you subtracted 2t from both sides: \[\large 2t-5 = t^2 -8t -9 \] add 5 to both sides \[\large 2t = t^2 -8t -4\] subtract 2t from both sides \[\large 0 = t^2 -10t -4\] Now use the quadratic formula as @ivancsc1996 showed (it won't factor)

OpenStudy (anonymous):

ohh I see... I added instead of subtracting... Okay thank you